# On the generalized eigenvalue problem in subspace-based excited state methods for quantum computers

### v1

@@ Abstract
Solving challenging problems in quantum chemistry is one of the leading promised applications of quantum computers. Within the quantum algorithms proposed for problems in excited state quantum chemistry, subspace-based quantum algorithms, including quantum subspace expansion (QSE), quantum equation of motion (qEOM) and quantum self-consistent equation-of-motion (q-sc-EOM), are promising for pre-fault-tolerant quantum devices. The working equation of QSE and qEOM requires solving a generalized eigenvalue equation with associated matrix elements measured on a quantum computer. Our quantitative analysis of the QSE method shows that the errors in eigenvalues increase drastically with an increase in the condition number of the overlap matrix when a generalized eigenvalue equation is solved in the presence of statistical sampling errors. This makes such methods unstable to errors that are unavoidable when using quantum computers. Further, at very high condition numbers of overlap matrix, the QSE’s working equation could not be solved without any additional steps in the presence of sampling errors as it becomes ill-conditioned. It was possible to use the thresholding technique in this case to solve the equation, but the solutions achieved had missing excited states, which may be a problem for future chemical studies. We also show that excited state methods that have an eigenvalue equation as the working equation, such as q-sc-EOM, do not have such problems and could be suitable candidates for excited-state quantum chemistry calculations using quantum computers.
@@ I Introduction
Quantum computing is expected to greatly enhance the simulation of quantum systems, particularly challenging problems in the ground state and excited state quantum chemistry  [1, 2, 3, 4, 5, 6, 7, 8, 5]. Recent years have seen rapid advances in methods and algorithms for computing the eigenstates of a molecular Hamiltonian in quantum chemistry using quantum computers  [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. Excited state quantum chemistry is exciting for quantum computing applications because classical computer-based quantum chemistry methods face challenges in several cases where an accurate description of excited states is needed, and quantum computers may provide an advance in these problems. For instance, leading scalable exited state methods face challenges in the accurate description of doubly excited states [20, 21, 22], accurate potential energy surfaces for the study of conical intersections and excited state dynamics [23, 24, 25, 26], especially when the ground state has a multireference character [27, 28].
@@
Two classes of quantum algorithms have been proposed to compute the potential energy surfaces of excited states: state-specific methods and subspace diagonalization-based methods. State-specific methods generally find one or a few excited states at a time by an optimization process similar to the ground state. Several methods in this type are proposed based on quantum phase estimation [29, 30, 31, 32, 33, 34, 35, 36, 37, 38], quantum annealing [39, 40] and variational quantum eigensolver (VQE) [41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53]. Although some of these methods may provide the exact excited state energies, they may be limited in their predictive power in near-degenerate excited states on imperfect quantum computers. For pre-fault-tolerant quantum computers, subspace diagonalization-based methods are promising (for a review of subspace methods for quantum computing applications, please refer to Motta et al. [54]). They provide a range of excited states by diagonalizing the Hamiltonian in a subspace created by electronic excitations on the ground state wavefunction. In these methods, a generalized eigenvalue equation or an eigenvalue equation is solved on a classical computer while the elements of these matrices are measured on a quantum computer. Leading examples of methods of this type are Quantum Subspace Expansion (QSE) [55, 56], Quantum Equation of Motion (qEOM) [57], and Quantum Self-consistent Equation-of-Motion (q-sc-EOM) [10, 58]. Several approaches for including the extended orbital space in subspace-based excited state methods have also been developed [11, 59, 60, 61, 62, 63]. Recently developed MORE-ADAPT by Grimsley et al.  [64] is also a related method, which optimises an adaptively built circuit for a weighted cost function with chosen initial states to provide accurate excited states. It requires solving an eigenvalue equation but also needs increased quantum resources. Another promising recent development in subspace methods is the sampled-based quantum diagonalization (SQD) [65, 66], which diagonalizes Hamiltonian in a subspace built using samples from a quantum state, requires solving an eigenvalue equation on a classical computer similar to q-sc-EOM.
@@
We have previously shown that q-sc-EOM has major theoretical and practical benefits compared with QSE and qEOM [10]. q-sc-EOM is size-intensive, unlike QSE and satisfies killer condition, unlike qEOM. The q-sc-EOM method, unlike QSE and qEOM, can be used with the Davidson algorithm developed by Kim et al. [67] to reduce shot requirements by an order of O(N4). Our qualitative tests have also shown that q-sc-EOM may be more robust to errors in measurement compared with QSE and qEOM [10]. This is due to the fact that the generalized eigenvalue problem, that is the working equation of QSE and qEOM, becomes sensitive to errors, even becoming singular when the condition number of the overlap matrix becomes too large. Singularities were reported in practical calculations using QSE by Barrison et al. in Ref. [68]. Singularities were also reported in the related quantum linear response formalisms that make use of generalized eigenvalue equations by Kjellgren et al. in Ref. [69]. They argued that the choice of orbitals becomes important in such cases. This problem is also seen in other ground state quantum algorithms which have similar generalized eigenvalue equations as the working equation, such as quantum Krylov-based methods [70, 71, 72, 73, 74], non-orthogonal quantum eigensolver [16, 75], connected moment expansion [76, 15], etc. Epperly et al. carried out a detailed study that lays down the mathematical foundations of such subspace diagonalization [77]. They show that the thresholding technique, which works by deleting small valued eigenvalues and associated eigenvectors of the overlap matrix and re-projecting the Hamiltonian on the resulting space, is effective at accurately solving for the lowest eigenvalue in generalized eigenvalue problems in case of singularities.
In this paper, we report a quantitative analysis of excited state methods, QSE and q-sc-EOM, that require solving a generalized eigenvalue equation and an eigenvalue equation, respectively, in the presence of statistical sampling errors. We show that solving the generalized eigenvalue equation makes the QSE method sensitive to errors in measurement. The problem is not just the presence of singularities (as reported previously), but the molecular calculations carried out using QSE become increasingly more error-prone as the condition number of the overlap matrix increases. In the case when the condition number becomes too large, the equation cannot be solved without the use of the thresholding technique [77]. Although thresholding provides a solution, it results in the loss of several eigenstates in the spectrum in the systems studied. Here, we report the effects of only sampling errors in this study; other noise sources in measurement on a quantum computer will also destabilize the QSE method (and similar excited state methods) due to similar arguments. We further show that q-sc-EOM is unaffected by the problems arising from the condition number of the overlap matrix. This is because the overlap matrix in q-sc-EOM is an identity matrix due to an orthonormal set of excited wavefunctions used.
@@ II.1 Quantum subspace expansion
QSE algorithm involves creating a subspace using the action of excitation operators on the ground state wavefunction (|$\ket{\Psi_{gr}}$) prepared on a quantum computer using any quantum algorithm. These excitations are used from an excitation manifold given by, $\{\hat{I}\cup\hat{G}_{I}\}$, where $\hat{G}_{I}$ includes single and double electronic excitations, given by
@@
and $\hat{I}$ is the Identity operator. Here, $\hat{a}^{\dagger}_{a}(\hat{a}_{a})$ are creation(destruction) operators acting on the orbital $a$. If |$\ket{\Psi_{VQE}}$ is the ground state wavefunction prepared using a VQE algorithm, the subspace can be created by the action of the excitation operators on |$\ket{\Psi_{VQE}}$.
@@
$\mathbf{H}_{I,J}$ $=\bra{\Psi_{\text{VQE}}}\hat{G}^{\dagger}_{I}\hat{H}\hat{G}_{J}\ket{\Psi_{\text{VQE}}},$
$\mathbf{S}_{I,J}$ $=\bra{\Psi_{\text{VQE}}}\hat{G}^{\dagger}_{I}\hat{G}_{J}\ket{\Psi_{\text{VQE}}},$
where $\hat{H}$ is the molecular Hamiltonian. Once the matrices are formed, the generalized eigenvalue problem is solved on a quantum computer, resulting in the energy of the ground and excited states of the quantum system as eigenvalues $E_{k}$.
@@ II.2 Quantum self-consistent equation of motion
The q-sc-EOM method is based on the equation of motion theory, which was developed for classical quantum chemistry by Rowe et al. [78] and first applied to quantum computing by Ollitrault et al. [57]. The q-sc-EOM method introduced self-consistent operators, originally developed by Prasad et al. [79], within the qEOM framework [10]. The working equation of the q-sc-EOM method takes a similar form as QSE. It involves solving the eigenvalue equation, given by
@@
$\mathbf{H}_{I,J}=\bra{\Psi_{\text{HF}}}\hat{G}_{I}^{\dagger}U^{\dagger}(\theta)\hat{H}U(\theta)\hat{G}_{J}\ket{\Psi_{\text{HF}}}-\delta_{I,J}E_{\text{HF}}.$ (3)
@@
|$\ket{\Psi_{\text{VQE}}}=U(\theta)\ket{\Psi_{\text{HF}}}$ (4)
and |$\ket{\Psi_{HF}}$ is the Hartree-Fock state. Finally, the simple eigenvalue equation is solved (or the matrix H is diagonalized) on a classical computer to find the energy eigenvalues of a molecular system.
@@ II.3 Eigenvalue problem vs generalized eigenvalue problem
A key difference between the two methods is that q-sc-EOM requires solving an eigenvalue problem since the overlap matrix is strictly an identity matrix proved analytically. QSE on the other hand, requires solving a generalized eigenvalue equation on a classical computer. In the presence of errors in the $\mathbf{H}$ and $\mathbf{S}$ matrix, this step can be challenging. This is because solving a generalized eigenvalue equation requires two steps, where the first step is the conversion of the generalized eigenvalue equation into an eigenvalue equation as
$(\mathbf{S}^{-1}\mathbf{H})C=CE,$ (5)
@@
which requires finding the inverse of the overlap matrix ($\mathbf{S}^{-1}$). The second step in this process is solving the new eigenvalue problem. Inverting a matrix in the first step in the presence of errors in the $\mathbf{S}$ matrix can be a challenging step, and its accuracy would depend strongly on the condition number of the overlap matrix and the error rates. If the condition number of the $\mathbf{S}$ matrix is large, it can become ill-conditioned.

### v2

@@ Abstract
Solving challenging problems in quantum chemistry is one of the most promising applications of quantum computers. Within the quantum algorithms proposed for problems in excited state quantum chemistry, subspace-based quantum algorithms, including quantum subspace expansion (QSE), quantum equation of motion (qEOM) and quantum self-consistent equation-of-motion (q-sc-EOM), are promising for pre-fault-tolerant quantum devices. The working equation of QSE and qEOM requires solving a generalized eigenvalue equation with associated matrix elements measured on a quantum computer. Our careful analytical and numerical analysis of the standard and generalized eigenvalue problems, especially in the context of excited-state methods, shows that the errors in eigenvalues magnify drastically with an increase in the condition number of the overlap matrix when a generalized eigenvalue equation is solved in the presence of statistical sampling errors. This makes methods such as QSE unstable to errors that are unavoidable when using quantum computers. Further, at very high condition numbers of the overlap matrix, the QSE’s working equation could not be solved without any additional steps in the presence of sampling errors, as it becomes ill-conditioned. It was possible to use the thresholding technique in this case to solve the equation, but the solutions achieved had missing excited states, which may be a problem for future chemical studies. We also show that excited-state methods that have an eigenvalue equation as the working equation, such as q-sc-EOM, do not have the problems associated with the condition number and could be generally more stable to errors, and therefore, more suitable candidates for excited-state quantum chemistry calculations using quantum computers.
@@ I Introduction
Quantum computing is expected to greatly enhance the simulation of quantum systems, particularly challenging problems in ground- and excited-state quantum chemistry [79, 47, 18, 58, 45, 32, 17, 23, 45]. Recent years have seen rapid advances in methods and algorithms for computing the eigenstates of a molecular Hamiltonian in quantum chemistry using quantum computers  [25, 4, 40, 9, 69, 70, 38, 30, 43, 5, 49]. Excited state quantum chemistry is exciting for quantum computing applications because classical computer-based quantum chemistry methods face challenges in several cases where an accurate description of excited states is needed, and quantum computers may provide an advance in these problems. For instance, leading scalable excited state methods face challenges in the accurate description of doubly excited states [15, 63, 61], accurate potential energy surfaces for the study of conical intersections and excited state dynamics [37, 78, 36, 76], especially when the ground state has a multi-reference character [44, 22].
@@
Two classes of quantum algorithms have been proposed to compute the potential energy surfaces of excited states: state-specific methods and subspace diagonalization-based methods. State-specific methods generally find one or a few excited states at a time by an optimization process similar to the ground state. Several methods in this type are proposed based on quantum phase estimation [34, 53, 65, 10, 73, 66, 55, 72, 71, 59], quantum annealing [77, 74] and variational quantum eigensolver (VQE) [19, 88, 87, 16, 68, 86, 50, 54, 80, 52, 84, 13, 85]. Although some of these methods may provide the exact excited state energies, they may be limited in their predictive power in near-degenerate excited states on imperfect quantum computers. For pre-fault-tolerant quantum computers, subspace diagonalization-based methods are promising (for a review of subspace methods for quantum computing applications, please refer to Motta et al. [51]). They provide a range of excited states by diagonalizing the Hamiltonian in a subspace created by electronic excitations on the ground state wavefunction. In these methods, a generalized eigenvalue equation or an eigenvalue equation is solved on a classical computer while the elements of these matrices are measured on a quantum computer. Leading examples of methods of this type are Quantum Subspace Expansion (QSE) [48, 82], Quantum Equation of Motion (qEOM) [56], and Quantum Self-consistent Equation-of-Motion (q-sc-EOM) [4, 39]. Several approaches for including extended orbital space in subspace-based excited state methods have also been developed [40, 31, 62, 83, 12, 11]. Recently developed MORE-ADAPT by Grimsley et al.  [26] is also a related method, which optimizes an adaptively built circuit for a weighted cost function with chosen initial states to provide accurate excited states. It requires solving an eigenvalue equation but also needs increased quantum resources. Another promising recent development in subspace methods is the sampled-based quantum diagonalization (SQD) [8, 67], which diagonalizes Hamiltonian in a subspace built using samples from a quantum state, requires solving an eigenvalue equation on a classical computer similar to q-sc-EOM.
@@
We have previously shown that q-sc-EOM has major theoretical and practical benefits compared with QSE and qEOM [4]. q-sc-EOM is size-intensive, unlike QSE and satisfies killer condition, unlike qEOM. The q-sc-EOM method, unlike QSE and qEOM, can be combined with the Davidson algorithm developed by Kim et al. [33] to reduce shot requirements by a factor of $O(N^{4})$. Generally applicable error mitigation strategies for the associated quantum linear response formalism were introduced in Ref. [91]. Our analytical and numerical tests have also shown that q-sc-EOM may be more robust to errors in measurement compared with QSE and qEOM [4]. This is due to the fact that the generalized eigenvalue problem, that is the working equation of QSE and qEOM, becomes sensitive to errors, even becoming singular when the condition number of the overlap matrix becomes too large. Singularities were reported in practical calculations using QSE by Barrison et al. in Ref. [7]. Singularities were also reported in the related quantum linear response formalisms that make use of generalized eigenvalue equations by Kjellgren et al. in Ref. [35]. They argued that the choice of orbitals becomes important in such cases. This problem is also seen in other ground state quantum algorithms which have similar generalized eigenvalue equations as the working equation, such as quantum Krylov-based methods [20, 89, 2, 81, 57], non-orthogonal quantum eigensolver [30, 6], connected moment expansion [90, 38], etc. Epperly et al. carried out a detailed study that lays down the mathematical foundations of such subspace diagonalization [21]. They show that the thresholding technique, which works by deleting small valued eigenvalues and associated eigenvectors of the overlap matrix and re-projecting the Hamiltonian on the resulting space, is effective at accurately solving for the lowest eigenvalue in generalized eigenvalue problems in case of singularities.
In this paper, we report both analytical and numerical analyses of excited-state methods, QSE and q-sc-EOM, that require solving a generalized eigenvalue equation and an eigenvalue equation, respectively, in the presence of statistical sampling errors. Our analysis results in the following contributions in this paper: (a) we translate perturbation bounds into practical diagnostics linking $\lambda_{\min}(\mathbf{S})$, expected shot noise, and excited-state reliability, (b) show numerically that ‘moderate’ condition numbers can already induce large instability under shot noise in QSE (H4 and NH3 cases), (c) demonstrate that thresholding can stabilize solvability but can reduce spectral completeness (missing excited states) in the studied examples, (d) contrast with an orthonormal-overlap formulation (q‑sc‑EOM) to isolate the role of generalized eigenproblem conditioning.
The sensitivity of generalized eigenvalue problems to ill-conditioning is a classical result in numerical linear algebra and has been recognized in prior quantum subspace diagonalization work (including Ref.  [21]). Our contribution is not a new perturbation theorem; rather, we quantify and illustrate how this mechanism manifests in realistic excited-state subspace algorithms with unavoidable sampling noise, and we highlight consequences that matter for chemistry use cases (e.g., spectral incompleteness under stabilization).
@@ II.1 Quantum subspace expansion
[…]
@@
[…]
@@
[…]
where $\hat{H}$ is the molecular Hamiltonian. Once the matrices are formed, the generalized eigenvalue problem is solved on a classical computer, resulting in the energy of the ground and excited states of the quantum system as eigenvalues $E_{k}$.
@@ II.2 Quantum self-consistent equation of motion
The q-sc-EOM method is based on the equation of motion theory, which was developed for classical quantum chemistry by Rowe et al. [64] and first applied to quantum computing by Ollitrault et al. [56]. The q-sc-EOM method introduced self-consistent operators, originally developed by Prasad et al. [60], within the qEOM framework [4]. The working equation of the q-sc-EOM method takes a similar form to QSE. It involves solving the eigenvalue equation, given by
@@
[…]
@@
[…]
@@ II.3 Eigenvalue problem vs generalized eigenvalue problem
A key difference between the two methods is that q-sc-EOM requires solving an eigenvalue problem since the overlap matrix is strictly an identity matrix proved analytically. QSE on the other hand, requires solving a generalized eigenvalue equation on a classical computer. In the presence of errors in the $\mathbf{H}$ and $\mathbf{S}$ matrix, this step can be challenging. This is because solving a generalized eigenvalue equation requires two steps, where the first step is the conversion of the generalized eigenvalue equation into an eigenvalue equation by modifying the eigenvector matrix as
$C=XC^{\prime},$ (5)
where $X=\mathbf{S}^{-\frac{1}{2}}$ and multiplying by $X^{\dagger}$ on the left. This results in an effective eigenvalue equation
$\mathbf{H}^{\prime}C^{\prime}=C^{\prime}E,$ (6)
where $\mathbf{H}^{\prime}=X^{\dagger}\mathbf{H}X$.
@@
This process requires finding the inverse of the overlap matrix ($\mathbf{S}^{-\frac{1}{2}}$). The second step in this process is solving the new eigenvalue problem. Inverting a matrix in the first step in the presence of errors in the $\mathbf{S}$ matrix can be a challenging step, and its accuracy would depend strongly on the condition number of the overlap matrix and the error rates. Conventional knowledge suggests that if the condition number of the $\mathbf{S}$ matrix is large, it can become unstable.
If the condition number is “too large”, it can lead to singularities as seen in many other studies on classical and quantum algorithms. Quantum algorithms are especially vulnerable because, while classical computers work at numerical accuracy of $\sim$10-16, even perfect quantum computers will always have statistical sampling errors that are inherent in quantum computation. In prior studies, many efforts have been made to solve the problem of singularities in generalized eigenvalue equations. Several practical methods have been employed for solving such problems, including the thresholding method [21], the deflation method [27, 28, 24, 46], and the shift method [27, 28, 24, 46], with thresholding as the most common choice for quantum algorithms. The thresholding technique involves re-projecting the Hamiltonian in a reduced subspace of eigenvectors of the overlap matrix, with corresponding eigenvalues greater than a threshold.
II.4 Error analysis in standard and generalized eigenvalue equation
In this subsection, we will provide a concise theoretical analysis of errors in eigenvalues in standard and generalized eigenvalue equations using matrix perturbation theory. For detailed analysis, please refer to Ref. [21].
II.4.1 Standard eigenvalue equation
In our study, the Hamiltonian, $\mathbf{H}$, is Hermitian and the overlap matrix, $\mathbf{S}$, is positive definite. The standard eigenvalue equation is given by
$\mathbf{H}x=\lambda x,$ (7)
where $\lambda$ and $x$ are the eigenvalue and eigenvector, respectively. The perturbed matrix and eigenvalues can be written as
$\tilde{\mathbf{H}}$ $=\mathbf{H}+\Delta\mathbf{H},$ (8)
$\tilde{\lambda}$ $=\lambda+\delta\lambda.$ (9)
Here, $\Delta\mathbf{H}_{i,j}\leq\epsilon$, with $\epsilon$ being the maximum allowed error (dependent on the number of shots used). The perturbed eigenvalue equation can be written as
$(\mathbf{H}+\Delta\mathbf{H})(x+\delta x)=(\lambda+\delta\lambda)(x+\delta x).$ (10)
Taking the first-order terms in error and rearranging, we get the first-order eigenvalue shift
$\delta\lambda^{(1)}=x^{\dagger}\Delta\mathbf{H}\,x.$ (11)
For any vector $x$, we know that [29]
[…]
where $||\cdot||_{2}$ and $||\cdot||_{F}$ are the spectral and Frobenius norms, respectively. Since,
[…]
where $n$ is the dimension of $\mathbf{H}$. We can state that
$|\delta\lambda^{(1)}|$ $\leq n\epsilon,$ (15)
$|\tilde{\lambda}-\lambda|$ $=O(n\epsilon)+O(\epsilon^{2}).$ (16)
This means that the error in eigenvalues for the standard eigenvalue equation scales linearly in $\epsilon$ with a worst-case pre-factor of $n$.
II.4.2 Generalized eigenvalue equation
We start with the generalized eigenvalue equation
$\mathbf{H}x=\lambda\mathbf{S}x.$ (17)
We choose eigenvectors normalized in the $\mathbf{S}$-inner product as $x^{\dagger}\mathbf{S}x=1$. The perturbed generalized eigenvalue equation, following a similar process as in the standard eigenvalue equation, can be written as
$(\mathbf{H}+\Delta\mathbf{H})(x+\delta x)=(\lambda+\delta\lambda)(\mathbf{S}+\Delta\mathbf{S})(x+\delta x).$ (18)
Collecting first-order terms in error and rearranging, we get
$\delta\lambda^{(1)}=x^{\dagger}(\Delta\mathbf{H}-\lambda\Delta\mathbf{S})x.$ (19)
Using the operator norm bound, we have
[…]
Since $\mathbf{S}\succ 0$, all eigenvalues of $\mathbf{S}$ are positive, and
[…]
giving the relationship
[…]
Also, by the triangle inequality [29] and Eq. (14),
[…]
Finally, using Eqs. (20), (22) and (23), we get
$|\delta\lambda|$ $\leq\frac{n(1+|\lambda|)}{\lambda_{\min}(\mathbf{S})}\epsilon,$ (24)
$|\tilde{\lambda}-\lambda|$ $=O\Big(\frac{n(1+|\lambda|)}{\lambda_{\min}(\mathbf{S})}\epsilon\Big)+O(\epsilon^{2}).$ (25)
This means that the error in eigenvalues for the generalized eigenvalue equation still scales linearly in $\epsilon$ but with a worst-case pre-factor of $n\frac{(1+|\lambda|)}{\lambda_{\min}(\mathbf{S})}$. Thus, the errors magnify in case of small $\lambda_{\min}(\mathbf{S})$ or large $\lambda$.
II.4.3 Practical shot-noise guidance
The conservative perturbation bound derived above can be converted into a practical guideline in the presence of statistical sampling noise (shot noise). If each matrix element entering $\mathbf{H}$ and $\mathbf{S}$ is estimated using $N_{\text{shots}}$ measurements, then the typical statistical estimation error scales as
$\epsilon\approx\frac{\sigma}{\sqrt{N_{\text{shots}}}},$ (26)
where $\sigma$ is an observable- and estimator-dependent constant set by the measurement variance (often $\sigma=O(1)$ for bounded observables). Substituting this scaling into the generalized-eigenvalue bound gives the worst-case estimate
$|\delta\lambda|\lesssim\frac{n(1+|\lambda|)}{\lambda_{\min}(\mathbf{S})}\,\frac{\sigma}{\sqrt{N_{\text{shots}}}}.$ (27)
Therefore, to achieve a target eigenvalue accuracy $|\delta\lambda|\leq\delta$, the required number of shots per matrix element scales as
$N_{\text{shots}}\gtrsim\left(\frac{\sigma\,n(1+|\lambda|)}{\lambda_{\min}(\mathbf{S})\,\delta}\right)^{2}.$ (28)
Equivalently, for a fixed shot budget $N_{\text{shots}}$ and tolerance $\delta$, a necessary (worst-case) condition for reliability is
$\lambda_{\min}(\mathbf{S})\gtrsim\frac{\sigma\,n(1+|\lambda|)}{\delta\sqrt{N_{\text{shots}}}}.$ (29)
These relations make explicit that the sampling cost grows quadratically as $\lambda_{\min}(\mathbf{S})$ decreases (and, in typical cases where $\lambda_{\max}(\mathbf{S})=O(1)$, grows approximately as $\kappa(\mathbf{S})^{2}$). This scaling is qualitatively consistent with the rapid growth of the mean error and variance in QSE as the overlap-matrix condition number increases in the medium-condition-number regime (Fig. 2). Moreover, for the case with $\kappa(\mathbf{S})=592.5$, the substantially larger shot requirements observed for QSE relative to q-sc-EOM to achieve comparable accuracy (Fig. 5) are in line with the presence of the amplification factor $1/\lambda_{\min}(\mathbf{S})$ in the generalized-eigenvalue bound.
This diagnostic also suggests when stabilization becomes necessary: when $\lambda_{\min}(\mathbf{S})$ becomes comparable to (or smaller than) the effective noise floor set by $\epsilon$, the generalized-eigenvalue solve can become unreliable, motivating procedures such as thresholding. However, such stabilization effectively reduces the working subspace and may compromise spectral completeness, consistent with the missing states observed when thresholding is required in the high-condition-number regime (Fig. 6). Finally, the reduction in QSE errors observed when the overlap matrix is made artificially exact (Fig. 4) supports the interpretation that noise in $\mathbf{S}$, and its subsequent inversion in the generalized-eigenvalue solve, is a key driver of the observed instability. In contrast, for formulations with an orthonormal basis where $\mathbf{S}=\mathbf{I}$ (e.g., q-sc-EOM), $\lambda_{\min}(\mathbf{S})=1$ and this particular amplification mechanism is absent, consistent with the stability observed across the corresponding examples (e.g., Fig. 3).